# Regression-Based Single-Point Zeroth-Order Optimization

### v1

@@ A.1 More on Assumption 2
We first prove the following result (under a very mild condition stated in (31)) showing that Assumption 2 holds with $C_{d}=1$ in the case $d=1$. Define $[N]\coloneq\{1,2,\cdots,N\}$.
@@
Suppose Assumption 1 holds. Suppose $d=1$. Suppose also that there exists some $i\in[m-1]$ such that
$|\hat{\bm{x}}_{t,i}-\hat{\bm{x}}|>0.$ (31)
Then, with $\bm{g}_{t}$ being defined as in (19), for any $t\geq m$, Assumption 2 holds with $C_{d}=1$, i.e.
$|\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})|\leq\frac{L}{2}\max_{i=1:m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|.$
@@
Note that $d=1$. Since we assume there exists some $i\in[m-1]$ such that $|\hat{\bm{x}}_{t,i}-\hat{\bm{x}}|>0$, we know that $S_{t}S_{t}^{\top}\in\mathbb{R}$ is invertible, satisfying the condition in Lemma 3. Thus, using Lemma 3, we note that the result is equivalent to showing that
$|(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}|\leq\frac{L}{2}\max_{i=1:m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|.$
Let $\epsilon_{t,i}$ denote the $i$-th entry of $\epsilon_{t}\in\mathbb{R}^{m-1}$. By $L$-smoothness of $f$, we observe that for any $i\in[m-1]$, we have
@@
$|\epsilon_{t,i}|=\ |f(\hat{\bm{x}}_{t-i})-(f(\hat{\bm{x}}_{t})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\rangle)|\leq\ \frac{L}{2}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|^{2}.$
Using the above result, in the case $d=1$, we observe that
$|(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}|=$ $\ \Big{|}\frac{1}{\sum_{i=1}^{m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|^{2}}\sum_{i=1}^{m-1}(\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t})\epsilon_{t,i}\Big{|}$
$\leq$ $\ \frac{1}{\sum_{i=1}^{m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|^{2}}\max_{i=1:m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|\sum_{i=1}^{m-1}|\epsilon_{t,i}|$
$\leq$ $\ \frac{1}{\sum_{i=1}^{m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|^{2}}\max_{i=1:m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|\sum_{i=1}^{m-1}\frac{L}{2}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|^{2}$
$\leq$ $\ \frac{L}{2}\max_{i=1:m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|,$
as desired. ∎
A.1.1 Technical Challenges of Proving Assumption 2 for General Cases with $d>1$
From Lemma 3, suppose that $S_{t}S_{t}^{\top}$ is invertible, proving Assumption 2 is equivalent to showing that
$\|(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}\|\leq\frac{C_{d}L}{2}\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|,$
for some $C_{d}>0$. A natural approach to proving the above inequality is to use the standard bound
$\|(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}\|\leq\|(S_{t}S_{t}^{\top})^{-1}\|_{2}\|S_{t}\epsilon_{t}\|.$
However, it is (i) unclear how we might theoretically bound $\|(S_{t}S_{t}^{\top})^{-1}\|_{2}$, and (ii), empirically, when $d<m=\mathcal{O}(d)$ regression samples are used, on the ridge regression, we observe that $\|(S_{t}S_{t}^{\top})^{-1}\|_{2}$ scales as [^4] We recall that the notation $\Omega(\cdot)$ is defined as follows: for any non-negative variable $n$, saying that a positive function $f(n)$ is $\Omega(g(n))$, where $g(n)$ is some positive funcion of $n$, means that there exists some absolute constant $C_{1}>0$ and a number $n_{0}$ such that for all $n\geq n_{0}$, we have $f(n)\geq C_{1}g(n)>0$.
$\|(S_{t}S_{t}^{\top})^{-1}\|_{2}=\Omega\left(\frac{d^{3}}{\sum_{i=1}^{m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|^{2}}\right).$
Combining this with the argument in Lemma 4 bounding $\|\epsilon_{t,i}\|$ as
@@ A.1.1 Technical Challenges of Proving Assumption 2 for General Cases with $d>1$
$|\epsilon_{t,i}|=\ |f(\hat{\bm{x}}_{t-i})-(f(\hat{\bm{x}}_{t})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\rangle)|\leq\ \frac{L}{2}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|^{2},$
it yields
$\|(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}\|\leq$ $\ \|(S_{t}S_{t}^{\top})^{-1}\|_{2}\|S_{t}\epsilon_{t}\|$
$\leq$ $\ \Omega\left(\frac{d^{3}}{\sum_{i=1}^{m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|^{2}}\right)\left\|\sum_{i=1}^{m-1}(\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t})\epsilon_{t,i}\right\|$
$\leq$ $\ \Omega\left(\frac{d^{3}}{\sum_{i=1}^{m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|^{2}}\right)\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|\sum_{i=1}^{m-1}\frac{L}{2}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|^{2}$
$\leq$ $\ \Omega\left(d^{3}\right)\frac{L}{2}\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|,$
yielding $C_{d}=\Omega(d^{3})$, which as we later demonstrate empirically in Appendix C, vastly overestimates the true size of $C_{d}$ that tends to scale as $\mathcal{O}(\sqrt{d})$. This suggests that the natural approach of estimating $C_{d}$ by first bounding $\|(S_{t}S_{t}^{\top})^{-1}\|_{2}$ does not yield tight bounds; additionally, it is in the first place also non-trivial to theoretically prove an upper bound for $\|(S_{t}S_{t}^{\top})^{-1}\|_{2}$.
Thus, we regard the task of accurately bounding $\|(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}\|$ as a very challenging theoretical problem which we leave to future work.
A.2 Proofs for Lemma 1, Theorem 1, and Theorem 2
Throughout the analysis, without loss of generality, we assume that $C_{d}\geq 1$. We also recall what we stated earlier, which is that we will study a variant of the L-RESZO algorithm with an adaptive choice of smoothing radius $\delta_{t}$ as (20),
$\delta_{t}:=\eta\|\bm{g}_{t-1}\|.$

### v2

@@ B.1 Theoretical justification for Assumption 2
We can further simplify the above result to an even simpler characterization.
@@
Consider the cumulative sum matrix $C\in\mathbb{R}^{m\times m}$, defined as
$C\;=\;\begin{pmatrix}1&0&0&\cdots&0\\ 1&1&0&\cdots&0\\ 1&1&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1&1&1&\cdots&1\end{pmatrix}$.
Define also the matrix
$D_{t}^{\top}:=\begin{bmatrix}(\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t})^{\top}\\ (\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1})^{\top}\\ \vdots\\ (\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2})^{\top}\\ \end{bmatrix}$,
as well as the following error vector
$\hat{\epsilon}_{t}:=\begin{bmatrix}f(\hat{\bm{x}}_{t-1})-(f(\hat{\bm{x}}_{t})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t}\rangle)\\ f(\hat{\bm{x}}_{t-2})-\left(f(\hat{\bm{x}}_{t-1})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1}\rangle\right)\\ \vdots\\ f(\hat{\bm{x}}_{t-m+1})-\left(f(\hat{\bm{x}}_{t-m+2})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2}\rangle\right)\end{bmatrix}$. (32)
Suppose $D$ has full rank. Then,
$\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})=(D_{t}D_{t}^{\top})^{-1}D_{t}\hat{\epsilon}_{t}.$
@@
From the earlier result, we saw that when $S$ is full rank (equivalent to $D$ being full rank), we have
$=$ $\ \nabla f(\hat{\bm{x}}_{t})+(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}.$
It can be easily verified that $S_{t}^{\top}=CD_{t}^{\top}$. Since $(S_{t}^{\top})^{\dagger}=(CD_{t}^{\top})^{\dagger}=(D_{t}^{\top})^{\dagger}C^{-1}$, we have that
$\bm{g}_{t}=$ $\ \nabla f(\hat{\bm{x}}_{t})+(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}$
$=$ $\ \nabla f(\hat{\bm{x}}_{t})+(D_{t}D_{t}^{\top})^{-1}D_{t}C^{-1}\epsilon_{t}.$
Since $C^{-1}$ is the following matrix
$C^{-1}\;=\;\begin{pmatrix}1&0&0&\cdots&0\\ -1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&0\\ 0&\cdots&0&-1&1\end{pmatrix}$,
the desired result follows since $C^{-1}\epsilon_{t}=\hat{\epsilon}_{t}.$
In order for Assumption 2 to hold, we need to rely on the following assumption on the smallest singular value of $D_{t}$. While difficult to prove since it involves computing the smallest singular value of random matrices with mild correlation across rows, we provide heuristic support for its validity.
Assumption 3.
Let $D_{t}^{\top}\in\mathbb{R}^{(m-1)\times d}$ be the matrix
$D_{t}^{\top}:=\begin{bmatrix}(\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t})^{\top}\\ (\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1})^{\top}\\ \vdots\\ (\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2})^{\top}\\ \end{bmatrix}$.
Then, the smallest singular value of $D_{t}^{\top}$ satisfies the following bound
$\sigma_{\min}(D_{t}^{\top})\geq\Omega\left(\frac{1}{\Gamma_{d,m}}\sqrt{\sum_{i=1}^{m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t-i+1}\|^{2}}\right).$
Here $\Gamma_{d,m}$ is a dimension-dependent constant satisfying
$\Gamma_{d,m}=\Theta\!\left(\sqrt{d}\sqrt{m}\right).$
Theoretical justification for Assumption 3.
We note that the matrix $D_{t}^{\top}$ can also be written as
$D_{t}^{\top}=$ $\ \begin{bmatrix}(\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t})^{\top}\\ (\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1})^{\top}\\ \vdots\\ (\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2})^{\top}\\ \end{bmatrix}$
$=$ $\ \begin{bmatrix}(\hat{\bm{x}}_{t-1}-{\bm{x}}_{t-1})^{\top}+({\bm{x}}_{t-1}-{\bm{x}}_{t})^{\top}+({\bm{x}}_{t}-\hat{\bm{x}}_{t})^{\top}\\ (\hat{\bm{x}}_{t-2}-{\bm{x}}_{t-2})^{\top}+({\bm{x}}_{t-2}-{\bm{x}}_{t-1})^{\top}+({\bm{x}}_{t-1}-\hat{\bm{x}}_{t-1})^{\top}\\ \vdots\\ (\hat{\bm{x}}_{t-m+1}-{\bm{x}}_{t-m+1})^{\top}+({\bm{x}}_{t-m+1}-{\bm{x}}_{t-m+2})^{\top}+({\bm{x}}_{t-m+2}-\hat{\bm{x}}_{t-m+2})^{\top}\\ \end{bmatrix}$
$=$ $\ \begin{bmatrix}(\bm{u}_{t-1})^{\top}+({\bm{x}}_{t-1}-{\bm{x}}_{t})^{\top}+(\bm{u}_{t})^{\top}\\ (\bm{u}_{t-2})^{\top}+({\bm{x}}_{t-2}-{\bm{x}}_{t-1})^{\top}+(\bm{u}_{t-1})^{\top}\\ \vdots\\ (\bm{u}_{t-m+1})^{\top}+({\bm{x}}_{t-m+1}-{\bm{x}}_{t-m+2})^{\top}+(\bm{u}_{t-m+2})^{\top}\\ \end{bmatrix}$
$=$ $\ G_{t}^{\top}+H_{t}^{\top},$
where
$G_{t}^{\top}=$ $\ \begin{bmatrix}({\bm{x}}_{t-1}-{\bm{x}}_{t})^{\top}\\ ({\bm{x}}_{t-2}-{\bm{x}}_{t-1})^{\top}\\ \vdots\\ ({\bm{x}}_{t-m+1}-{\bm{x}}_{t-m+2})^{\top}\end{bmatrix},\quad H_{t}^{\top}=\begin{bmatrix}(\bm{u}_{t-1})^{\top}-(\bm{u}_{t})^{\top}\\ (\bm{u}_{t-2})^{\top}-(\bm{u}_{t-1})^{\top}\\ \vdots\\ (\bm{u}_{t-m+1})^{\top}-(\bm{u}_{t-m+2})^{\top}\\ \end{bmatrix}$,
with $\bm{u_{j}}^{\top}$ is sampled independently and uniformly from the sphere with norm $\|\bm{x}_{j}-\bm{x}_{j-1}\|$. To find the smallest singular value of $D_{t}^{\top}$, we thus have to find the smallest eigenvalue of the matrix $(D_{t}D_{t}^{\top})=(G_{t}G_{t}^{\top}+G_{t}H_{t}^{\top}+H_{t}G_{t}^{\top}+H_{t}H_{t}^{\top})$. While $G_{t}$ and $H_{t}$ are slightly correlated, the cross-terms should display concentration effects and thus dominated by the effects of the terms $G_{t}G_{t}^{\top}+H_{t}H_{t}^{\top}$. To see this, taking $G_{t}H_{t}^{\top}$ for example, we observe that
$G_{t}H_{t}^{\top}=\sum_{j=1}^{m-1}(\bm{x}_{t-j}-\bm{x}_{t-j+1})(\bm{u}_{t-j}-\bm{u}_{t-j+1})^{\top}.$ (33)
Since $\bm{u}_{t-j+1}$ is independent of $\bm{x}_{t-j}-\bm{x}_{t-j+1}$, the term $\sum_{j=1}^{m-1}(\bm{x}_{t-j}-\bm{x}_{t-j+1})(\bm{u}_{t-j+1})^{\top}$ is mean zero and exhibits concentration of measure effects. Similarly, any correlation between $\bm{x}_{t-j}-\bm{x}_{t-j+1}$ and $\bm{u}_{t-j}$ is likely mild; when the iterates are generated before the regression starts, e.g., via one-point method, $\bm{x}_{t-j}-\bm{x}_{t-j+1}$ will have no relation to $\bm{u}_{t-j}$, since the random direction at iterate tjt-j is independent of the perturbation $\bm{u}_{t-j}$ which is only to obtain the function value evaluation. When $\bm{x}_{t-j}-\bm{x}_{t-j+1}=\eta\bm{g}_{t-j}$ is the result of a regression-based gradient estimation update, $\bm{g}_{t-j}$ is affected by $\bm{u}_{t-j}$, but only in one row, and thus unlikely to have significant correlation.
We may thus assume that
$\sigma_{\min}(D_{t}D_{t}^{\top})=\Theta(\sigma_{\min}(G_{t}G_{t}^{\top}+H_{t}H_{t}^{\top}))\geq\Theta(\sigma_{\min}(H_{t}H_{t}^{\top})).$
Fully characterizing the distribution of $\sigma_{\min}(H_{t}H_{t}^{\top})$ is highly non-trivial since the norm of each row of $H_{t}^{\top}$ has a different variance, namely that the $j$-th row of $H_{t}^{\top}$, which is equal to $(\bm{u}_{t-j}-\bm{u}_{t-j+1})^{\top}$, has variance $\|\bm{u}_{t-j}\|^{2}+\|\bm{u}_{t-j+1}\|^{2}=\|\bm{x}_{t-j}-\bm{x}_{t-j-1}\|^{2}+\|\bm{x}_{t-j+1}-\bm{x}_{t-j}\|^{2}$.
To proceed, note that $H_{t}^{\top}\in\mathbb{R}^{(m-1)\times d}$ has rows $h_{j}^{\top}:=(\bm{u}_{t-j}-\bm{u}_{t-j+1})^{\top}$ for $j=1,\dots,m-1$. To build intuition for the scale of $\sigma_{\min}(H_{t}^{\top})$, consider the Gram matrix
$H_{t}H_{t}^{\top}\;=\;\sum_{j=1}^{m-1}h_{j}h_{j}^{\top}\;=\;\sum_{j=1}^{m-1}(\bm{u}_{t-j}-\bm{u}_{t-j+1})(\bm{u}_{t-j}-\bm{u}_{t-j+1})^{\top}.$
Expanding the summand gives
$(\bm{u}_{t-j}-\bm{u}_{t-j+1})(\bm{u}_{t-j}-\bm{u}_{t-j+1})^{\top}=\bm{u}_{t-j}\bm{u}_{t-j}^{\top}+\bm{u}_{t-j+1}\bm{u}_{t-j+1}^{\top}-\bm{u}_{t-j}\bm{u}_{t-j+1}^{\top}-\bm{u}_{t-j+1}\bm{u}_{t-j}^{\top}.$
Summing over $j$ yields
$H_{t}H_{t}^{\top}=\underbrace{\sum_{j=1}^{m-1}\bm{u}_{t-j}\bm{u}_{t-j}^{\top}+\sum_{j=1}^{m-1}\bm{u}_{t-j+1}\bm{u}_{t-j+1}^{\top}}_{\text{``diagonal'' terms}}\;-\;\underbrace{\sum_{j=1}^{m-1}\big(\bm{u}_{t-j}\bm{u}_{t-j+1}^{\top}+\bm{u}_{t-j+1}\bm{u}_{t-j}^{\top}\big)}_{\text{cross terms}}.$
The diagonal part is essentially a weighted sum of rank-one matrices of the form $\bm{u}_{k}\bm{u}_{k}^{\top}$. The cross terms are mean-zero (conditionally on $\{\bm{x}_{k}\}$) because the directions $\bm{u}_{k}$ are independent across $k$ and isotropic, so $\mathbb{E}[\bm{u}_{t-j}\bm{u}_{t-j+1}^{\top}]=0$. Consequently, by concentration of measure, the cross terms exhibit cancellations and are typically of smaller spectral norm compared to the diagonal terms. This motivates the approximation
$\sigma_{\min}(H_{t}H_{t}^{\top})\;\approx\;\sigma_{\min}\!\left(\sum_{k=t-m+2}^{t-1}\bm{u}_{k}\bm{u}_{k}^{\top}\right).$
Next, assume that $\|\bm{u}_{k}\|^{2}$ are comparable over the window, i.e., $\|\bm{u}_{k}\|^{2}\asymp r^{2}$ for $k=t-m+2,\dots,t-1$. Under either the Gaussian model $\bm{u}_{k}\sim\mathcal{N}(0,\tfrac{r^{2}}{d}I_{d})$ or the uniform-sphere model (with comparable sub-Gaussian tails), the matrix with rows proportional to $\bm{u}_{k}^{\top}$ is a (sub-)Gaussian design with unit-variance entries up to scaling. Therefore, by classic results on the smallest singular value of sub-Gaussian random matrices (e.g., Theorem 1.1 in [21]), we have heuristically that
$\sigma_{\min}\!\left(\sum_{k=t-m+2}^{t-1}\bm{u}_{k}\bm{u}_{k}^{\top}\right)\;=\;\Theta\!\left(\frac{r^{2}}{d}\right),$
which implies
$\sigma_{\min}(H_{t}^{\top})\;=\;\sqrt{\sigma_{\min}(H_{t}H_{t}^{\top})}\;\approx\;\Theta\!\left(\frac{r}{\sqrt{d}}\right).$
Finally, using $\|\bm{u}_{k}\|^{2}\asymp r^{2}$ and m1m-1 terms in the window,
$r\;\asymp\;\sqrt{\frac{1}{m-1}\sum_{k=t-m+2}^{t-1}\|\bm{u}_{k}\|^{2}},$
so the above scaling can be rewritten as
$\sigma_{\min}(H_{t}^{\top})\;\approx\;$ $\ \Theta\!\left(\frac{1}{\sqrt{d}}\sqrt{\frac{1}{m-1}\sum_{k=t-m+2}^{t-1}\|\bm{u}_{k}\|^{2}}\right)\;=\;\Theta\!\left(\frac{1}{\sqrt{d(m-1)}}\sqrt{\sum_{k=t-m+2}^{t-1}\|\bm{u}_{k}\|^{2}}\right)$
$=$ $\ \Omega\left(\frac{1}{\sqrt{d}\sqrt{m}}\sqrt{\sum_{i=1}^{m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t-i+1}\|^{2}}\right).$
We are now ready to prove Assumption 2, assuming that Assumption 3 holds.
Proof of Assumption 2.
Suppose Assumption 3 holds. Recall from Lemma 4 that
$\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})=(D_{t}D_{t}^{\top})^{-1}D_{t}\hat{\epsilon}_{t},$
where
$D_{t}^{\top}:=\begin{bmatrix}(\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t})^{\top}\\ (\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1})^{\top}\\ \vdots\\ (\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2})^{\top}\\ \end{bmatrix}$,
and
@@
$\hat{\epsilon}_{t}:=\begin{bmatrix}f(\hat{\bm{x}}_{t-1})-(f(\hat{\bm{x}}_{t})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t}\rangle)\\ f(\hat{\bm{x}}_{t-2})-\left(f(\hat{\bm{x}}_{t-1})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1}\rangle\right)\\ \vdots\\ f(\hat{\bm{x}}_{t-m+1})-\left(f(\hat{\bm{x}}_{t-m+2})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2}\rangle\right)\end{bmatrix}$.
From Assumption 3, we have that
$\sigma_{\min}(D_{t}^{\top})\geq\Omega\left(\frac{1}{\Gamma_{d,m}}\sqrt{\sum_{i=1}^{m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t-i+1}\|^{2}}\right).$
It follows then that
$\|\bm{g}_{t}-\nabla f(\hat{x}_{t})\|\leq\frac{1}{\sigma_{\min}(D_{t})}\|\hat{\epsilon}_{t}\|\leq\frac{\mathcal{O}(\sqrt{d}\sqrt{m})}{\sqrt{\sum_{i=1}^{m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t-i+1}\|^{2}}}\|\hat{\epsilon}_{t}\|.$
Using Taylor expansion, note we can also write $\hat{\epsilon}_{t}$ as
@@
$\hat{\epsilon}_{t}:=$ $\ \begin{bmatrix}f(\hat{\bm{x}}_{t-1})-(f(\hat{\bm{x}}_{t})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t}\rangle)\\ f(\hat{\bm{x}}_{t-2})-\left(f(\hat{\bm{x}}_{t-1})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1}\rangle\right)\\ \vdots\\ f(\hat{\bm{x}}_{t-m+1})-\left(f(\hat{\bm{x}}_{t-m+2})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2}\rangle\right)\\ \end{bmatrix}$
$=$ $\ \begin{bmatrix}\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t}\rangle+\mathcal{O}(\frac{L}{2})\|\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t}\|^{2}-(\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t}\rangle)\\ \langle\nabla f(\hat{\bm{x}}_{t-1}),\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1}\rangle+\mathcal{O}(\frac{L}{2})\|\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1}\|^{2}-(\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1}\rangle)\\ \vdots\\ \langle\nabla f(\hat{\bm{x}}_{t-m+2}),\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2}\rangle+\mathcal{O}(\frac{L}{2})\|\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2}\|^{2}-(\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2}\rangle)\\ \end{bmatrix}$
$=$ $\ \begin{bmatrix}\langle\nabla f(\hat{\bm{x}}_{t})-\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t}\rangle+\mathcal{O}(\frac{L}{2})\|\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t}\|^{2}\\ \langle\nabla f(\hat{\bm{x}}_{t-1})-\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1}\rangle+\mathcal{O}(\frac{L}{2})\|\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1}\|^{2}\\ \vdots\\ \langle\nabla f(\hat{\bm{x}}_{t-m+2})-\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2}\rangle+\mathcal{O}(\frac{L}{2})\|\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2}\|^{2}\\ \end{bmatrix}$
$=$ $\ \begin{bmatrix}\mathcal{O}(\frac{L}{2})\|\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t}\|^{2}\\ \mathcal{O}(L\|\hat{\bm{x}}_{t-1})-\nabla f(\hat{\bm{x}}_{t})\|\|\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1}\|)+\mathcal{O}(\frac{L}{2})\|\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1}\|^{2}\\ \vdots\\ \mathcal{O}(L\|\hat{\bm{x}}_{t-m+1})-\nabla f(\hat{\bm{x}}_{t})\|\|\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2}\|)+\mathcal{O}(\frac{L}{2})\|\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2}\|^{2}\\ \end{bmatrix}$.
Thus it follows that
$\|\hat{\epsilon}_{t}\|\leq\max\{\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|,\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t-i+1}\|\}\sqrt{\sum_{i=1}^{m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t-i+1}\|^{2}}$
It follows then that
$\|\bm{g}_{t}-\nabla f(\hat{x}_{t})\|\leq$ $\ \frac{1}{\sigma_{\min}(D_{t})}\|\hat{\epsilon}_{t}\|\leq\frac{\mathcal{O}(\sqrt{d}\sqrt{m})}{\sqrt{\sum_{i=1}^{m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t-i+1}\|^{2}}}\|\hat{\epsilon}_{t}\|$
$\leq$ $\ \mathcal{O}(\sqrt{d}\sqrt{m})\max\{\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|,\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t-i+1}\|\},$
i.e. the relation $C_{d,m}=\mathcal{O}(\sqrt{d}\sqrt{m})$ holds. [^2] We note that this is slightly different from what we stated in Assumption 2, but this result is identical to that in Assumption 2 for the purposes of our proofs.
B.2 Proofs for Lemma 1, Theorem 1, and Theorem 2
Throughout the analysis, without loss of generality, we denote $C_{d}:=C_{d,m}$ and assume that $C_{d}\geq 1$.